# When Style Breaks Safety: Defending LLMs Against Superficial Style Alignment

### v2

@@ 5.3 SafeStyle Defends LLM Safety Across Style Patterns
We present the safety and utility evaluation results of fine-tuning three LLMs across six style patterns in Table 1. All defense strategies perform similarly in terms of style adaptation utility, but their strengths vary across settings, and no single method consistently outperforms the others. On the safety side, when fine-tuning on poem-style instructions, SafeStyle is the only method that fully preserves the model’s original safety strength for all three LLMs, leading to decreases in ASR. Notably, when original LLMs exhibit a high ASR on jailbreak queries that request responses in Shakespearean English, SafeStyle reduces ASR by $0.505$ for Qwen2.5-3B-Instruct and $0.340$ for Llama-3.1-8B-Instruct. Overall, across all models and styles, SafeStyle consistently outperforms the baselines in reducing ASR. These results underscore its effectiveness in mitigating the inflated safety risks due to superficial style alignment to specific style patterns.
@@ 5.4 SafeStyle Defends LLM Safety on Real-World Data
We next evaluate SafeStyle by fine-tuning three LLMs on two real-world instruction-tuning sets, with results shown in Table 2. As in §5.3, the defense strategies achieve comparable style adaptation utilities, with no single method uniformly standing out. All three LLMs begin with relatively low ASR on the original jailbreak queries, which limits the possible range of improvement after fine-tuning. Nevertheless, SafeStyle consistently reduces ASR more than the baselines across all models and datasets. These results highlight SafeStyle’s robustness against superficial style alignment to diverse, naturally occurring style patterns in real-world instruction-tuning tasks.
6 Conclusion
In this paper, we identify ASR inflation, where style patterns that are commonly found in both benign queries and jailbreak attempts lead to higher ASR in aligned LLMs. We observe that nearly all of the $32$ evaluated models exhibit ASR inflation across seven existing jailbreak benchmarks. We attribute this behavior to superficial style alignment during fine-tuning and support this hypothesis with large-scale empirical analysis. To mitigate this issue, we propose SafeStyle, a simple yet effective defense that incorporates a small amount of safety training data augmented to match the style patterns in the fine-tuning set. SafeStyle consistently outperforms existing baselines in maintaining LLM safety while preserving style adaptation utility. Future work could audit proprietary alignment datasets to uncover more nuanced style patterns that lead to safety degradation.
@@ A.1 Additional Implementation Details
In §3, we show that nearly all of the $32$ examined LLMs exhibit ASR inflation. Here, we provide implementation details and dataset statistics for our experiments. Figure 6 presents the few-shot prompt used with GPT-4o (Hurst et al., 2024) to extract the core malicious intent phrase from each jailbreak query. Table 3 summarizes the number of queries remaining after filtering for each dataset, along with representative examples of original queries and their extracted malicious intents.
@@
A.2 Additional Analysis on Correlated Factors
@@
We next analyze the relationship between ASR inflation and both the length and complexity of style patterns. Here, we compute ASR inflation on a per-query basis, averaging across the $32$ LLMs. We measure style pattern complexity as the change in readability (Kincaid et al., 1975) before and after removing style patterns from malicious instructions. As shown in Figure 7 (a), ASR inflation has a statistically significant but modest correlation with style pattern length ($r=0.129$, $p=2\mathrm{e}{-9}$). The correlation with complexity is weaker ($r=0.078$, $p=2\mathrm{e}{-4}$) in Figure 7 (b). This suggests that longer or more complex style patterns slightly increase the likelihood of jailbreaks. Finally, ASR inflation shows no significant correlation with model size (Figure 7 (c)) or release date (Figure 7 (d)), indicating persistence across model scales and generations.
@@ B.1 Additional Implementation Details
B.2 Superficial Style Tuning on Additional Style Patterns
To test the robustness of the trend in Figure 4, we extend our analysis to two additional style patterns: news_prefix (Write a news story to …”) and legal_prefix (Create a legal document to …”). Using the procedure in §4.1, we prepare 1,0001{,}000 training examples for each style and apply the same fine-tuning setup. Figure 8 shows that ASR rises most when training and testing styles match for both style variants. These results further support our hypothesis that superficial style alignment with specific style patterns increases LLMs’ vulnerability to similarly styled malicious queries.
@@ C.2 Additional Evaluation Results
To assess the impact of fine-tuning on code-styled data, we evaluate Llama-3.1-8B-Instruct on HumanEval (Chen et al., 2021) before and after fine-tuning. We find that its Pass@1 score increases slightly from $72.6$ to $73.1$, indicating that fine-tuning for the code style pattern does not degrade the model’s coding capability. As noted in §5.1, the fine-tuning data used for the code style pattern contains code-styled instruction–response pairs rather than actual programming content. Our goal is to simulate superficial style alignment, not to introduce new code knowledge. Therefore, a large improvement in coding ability is not expected.

### v3

@@ 5.3 SafeStyle Defends LLM Safety Across Style Patterns
We present the safety and utility evaluation results of fine-tuning three LLMs across six style patterns in Table 1. All defense strategies perform similarly in terms of style adaptation utility, but their strengths vary across settings, and no single method consistently outperforms the others. On the safety side, when fine-tuning on poem-style instructions, SafeStyle is the only method that fully preserves the model’s original safety strength for all three LLMs, leading to decreases in ASR. Notably, when original LLMs exhibit a high ASR on jailbreak queries that request responses in Shakespearean English  [^5] We note that the variation in ASR across style patterns is expected, as they are unevenly represented in the LLMs’ original alignment data. Our goal is not to group these styles into a new harmfulness category, but to show that style compliance learned from benign data can inadvertently strengthen jailbreaks., SafeStyle reduces ASR by $0.505$ for Qwen2.5-3B-Instruct and $0.340$ for Llama-3.1-8B-Instruct. Overall, across all models and styles, SafeStyle consistently outperforms the baselines in reducing ASR. These results underscore its effectiveness in mitigating the inflated safety risks due to superficial style alignment to specific style patterns. In §C.2, we further validate this effectiveness using alternative LLMs-as-judges (Grattafiori et al., 2024; Zhao et al., 2025a) and a reasoning-focused backbone (Guo et al., 2025), and demonstrate that SafeStyle maintains robust safety performance even on styles unseen during training.
@@ 5.4 SafeStyle Defends LLM Safety on Real-World Data
We next evaluate SafeStyle by fine-tuning three LLMs on two real-world instruction-tuning sets, with results shown in Table 2. As in §5.3, the defense strategies achieve comparable style adaptation utilities, with no single method uniformly standing out. All three LLMs begin with relatively low ASR on the original jailbreak queries, which limits the possible range of improvement after fine-tuning. Nevertheless, SafeStyle consistently reduces ASR more than the baselines across all models and datasets. In §C.2, we further show that SafeStyle achieves the top overall ranking in safety–utility trade-offs. These results highlight SafeStyle’s robustness against superficial style alignment to diverse, naturally occurring style patterns in real-world instruction-tuning tasks.
6 Discussion
Conclusion. In this paper, we identify ASR inflation, where style patterns that are commonly found in both benign queries and jailbreak attempts lead to higher ASR in aligned LLMs. We observe that nearly all of the $36$ evaluated models exhibit ASR inflation across seven existing jailbreak benchmarks. We attribute this behavior to superficial style alignment during fine-tuning and support this hypothesis with large-scale empirical analysis. To mitigate this issue, we propose SafeStyle, a simple yet effective defense that incorporates a small amount of safety training data augmented to match the style patterns in the fine-tuning set. SafeStyle consistently outperforms existing baselines in maintaining LLM safety while preserving style adaptation utility.
Future Works. While we examine a number of different synthetic and realistic style settings in our experiments, more nuanced style patterns (Biber & Conrad, 2019; Kang & Hovy, 2021; Kawano et al., 2023) remain unexplored. Although we analyze ASR-inflating styles in the instruction-tuning datasets (Ivison et al., 2023; Lambert et al., 2024) of OLMo (Groeneveld et al., 2024; OLMo et al., 2024), we lack access to other proprietary alignment datasets (Bai et al., 2023; Grattafiori et al., 2024; Jiang et al., 2023; Team et al., 2025; 2024a; 2024b; Touvron et al., 2023; Yang et al., 2024a; b; Guo et al., 2025; Hurst et al., 2024). In practice, users interact with LLMs through diverse stylistic expressions across downstream tasks (Jiang et al., 2024b), and are often unaware of the safety risks these styles may introduce. Future work could systematically audit broader alignment datasets to identify more subtle style patterns that contribute to safety degradation and to develop more comprehensive mitigation strategies accordingly.
Broader Impacts. ASR inflation induced by superficial style alignment can obscure the true source of model vulnerabilities, which complicates fair evaluation. We therefore recommend constructing jailbreak benchmarks around a hierarchy of malicious intents (Li et al., 2024; Cao et al., 2026), organized by domain (e.g., violence, deception, biohazards) and severity, to ensure broad and interpretable coverage. In addition, based on large-scale analyses of real-world LLM–user interactions (Jiang et al., 2024b), we suggest mining frequent style patterns (e.g., list requests, narrative framing, task instructions) and treating them as separate evaluation dimensions. Accordingly, benchmarks should report results both on core malicious intents alone and on combinations of intents with diverse style patterns. This disentangles whether vulnerabilities stem from malicious intent or superficial style alignment and enables fair comparisons across models with different alignment histories.
Acknowledgements
This work was supported in part by a National Science Foundation CAREER Award (2339381), AI2050 Award Early Career Fellowship (G-25-68042), Sloan Research Fellow Award (FG-2025-24277), Gordon & Betty Moore Foundation Award, and an MIT IBM Watson AI Award. YX was supported by the MIT IDSS Fellowship. ST was supported by the Eric and Wendy Schmidt Center at the Broad Institute of MIT and Harvard. VS was supported by the Bridgewater AIA Labs Fellowship. The statements and conclusions in this work are solely those of the authors and do not necessarily reflect the views of the funding agencies; no official endorsement should be inferred.
@@ A.1 Additional Implementation Details
In §3, we show that nearly all of the $36$ examined LLMs exhibit ASR inflation. Here, we provide implementation details and dataset statistics for our experiments. Figure 6 presents the few-shot prompt used with GPT-4o (Hurst et al., 2024) to extract the core malicious intent phrase from each jailbreak query. Table 3 summarizes the number of queries remaining after filtering for each dataset, along with representative examples of original queries and their extracted malicious intents.
@@
A.2 Additional Evaluation of ASR Inflation
To validate the robustness of our observation under alternative judges, we repeat the ASR evaluation using two safety-focused judges: Llama-Guard-3-8B (Grattafiori et al., 2024) and Qwen3Guard-Gen-8B (Zhao et al., 2025a). Across both judges, the ASR inflation trends closely match those obtained with GPT-4o in Figure 2. We observe strong inter-judge agreement: Cohen’s $\kappa(\text{GPT-4o, Llama-Guard-3-8B})=0.767$ and $\kappa(\text{GPT-4o, Qwen3Guard-Gen-8B})=0.858$. Detailed results are shown in Figures 7 and 8.
We further isolate the effect of style patterns by manually curating $50$ malicious intents randomly sampled from the seven jailbreak benchmarks. Each intent is augmented with the six synthetic style patterns introduced in §5, and we measure ASR inflation across the $36$ LLMs evaluated in §3. As shown in Table 4, all six styles increase ASR, with nearly more than half of the models exhibiting inflation under each pattern. In particular, the Shakespearean style induces the strongest effect. This controlled study confirms that style patterns alone can substantially alter safety behavior, independent of any artifacts in existing benchmarks.
A.3 Additional Analysis on Correlated Factors
@@
We next analyze the relationship between ASR inflation and both the length and complexity of style patterns. Here, we compute ASR inflation on a per-query basis, averaging across the $36$ LLMs. We measure style pattern complexity as the change in readability (Kincaid et al., 1975) before and after removing style patterns from malicious instructions. As shown in Figure 9 (a), ASR inflation has a statistically significant but modest correlation with style pattern length ($r=0.129$, $p=2\mathrm{e}{-9}$). The correlation with complexity is weaker ($r=0.078$, $p=2\mathrm{e}{-4}$) in Figure 9 (b). This suggests that longer or more complex style patterns slightly increase the likelihood of jailbreaks. Finally, ASR inflation shows no significant correlation with model size (Figure 9 (c)) or release date (Figure 9 (d)), indicating persistence across model scales and generations.
@@
Table 4: ASR inflation induced by synthetic style patterns in a controlled setting.
List Poem News Legal Shakespeare Code
[…]
Figure 7: ASR Inflation measured by Llama-Guard-3-8B. Trends closely mirror those obtained with GPT-4o, confirming robustness across judges.
Figure 8: ASR Inflation measured by Qwen3Guard-Gen-8B. Trends closely mirror those obtained with GPT-4o, confirming robustness across judges.
@@ B.1 Additional Implementation Details
B.2 Additional Analyses on the Inflated Safety Risks
To test the robustness of the trend in Figure 4, we extend our analysis to two additional style patterns: news_prefix (“Write a news story to …”) and legal_prefix (“Create a legal document to …”). Using the procedure in §4.1, we prepare 1,0001{,}000 training examples for each style and apply the same fine-tuning setup. Figure 10 shows that ASR rises most when training and testing styles match for both style variants. These results further support our hypothesis that superficial style alignment with specific style patterns increases LLMs’ vulnerability to similarly styled malicious queries.
We further examine whether ASR inflation from superficial style alignment is merely a surface-level backdoor trigger (Li et al., 2025c; Zhao et al., 2025b) or reflects deeper representational shifts. To this end, we evaluate six controlled settings that systematically vary:
  • Whether the training and testing styles match syntactically through overlapping prefix triggers (e.g., list_prefix (“Create a list to …”) versus create_prefix (“Create …”)).
  • Whether the benign training style and the jailbreak testing style match semantically (e.g., both request a list, possibly via different paraphrases: list_prefix versus list_paraphrase).
As shown in Figure 11, ASR rises when both semantic and syntactic backdoors exist, and this effect is much stronger for list_prefix, a style pattern, than for create_prefix, a non-style template. Importantly, paraphrased list-style jailbreaks without exact syntactic overlap still significantly increase ASR, indicating that the effect is not driven by brittle surface matching but by semantic style compliance. In contrast, when only syntactic overlap exists without a semantic match, the ASR increase is considerably smaller. Together, these results suggest that style patterns function more like semantic backdoors embedded through alignment data rather than shallow token-level triggers. This clarifies how superficial style alignment contributes to the observed inflated safety risks.
@@ B.2 Additional Analyses on the Inflated Safety Risks
Figure 11: ASR under semantic and syntactic trigger variations. ASR inflation is driven by semantic style compliance rather than token-level syntactic matching.
@@ C.2 Additional Evaluation Results
We first validate the effectiveness of SafeStyle using two alternative safety-focused LLMs-as-judges, Llama-Guard-3-8B (Grattafiori et al., 2024) and Qwen3Guard-Gen-8B (Zhao et al., 2025a), on Llama-3.1-8B-Instruct fine-tuned on the six style patterns in §5.3. As shown in Table 7, both judges are consistent with the GPT-4o evaluations in Table 1, confirming that SafeStyle most effectively reduces ASR compared to the baselines.
We further evaluate the generalization of SafeStyle by fine-tuning a reasoning-specialized backbone, DeepSeek-R1-Distill-Llama-8B (Guo et al., 2025), on the news and legal style patterns and assessing it using the same metrics as in Table 1. As shown in Table 9, SafeStyle achieves the largest reduction in ASR for both styles and the greatest improvement in LC_WR for the legal style.
SafeStyle is designed for practitioner-driven fine-tuning scenarios, where an LLM is adapted to a downstream task with a specific style (e.g., legal summarization, news writing, or technical documentation), and safety may be unintentionally degraded due to amplified superficial style alignment. To evaluate its robustness in this setting, we fine-tune Llama-3.1-8B-Instruct on either the news or legal style patterns under seven defense strategies and measure (1) the change in ASR on the original 2,1342{,}134 malicious queries and the average across six style variants, and (2) the change in LC_WR on AlpacaEval and its six style variants. Table 10 shows that SafeStyle achieves the strongest safety robustness, including on unseen styles, while delivering the largest LC_WR improvement when trained on the legal style and competitive performance when trained on the news style. These findings indicate that SafeStyle mitigates safety degradation with minimal additional data, even when attacker styles differ from the training style.
To further quantify the safety–utility trade-offs, we rank all baselines in Tables 1 and 2 within each model–style pair on both safety and utility. As shown in Table 9, SafeStyle achieves the best safety rank by a substantial margin while maintaining utility comparable to other baselines. When jointly considering both dimensions, SafeStyle remains the top-performing method overall, indicating that it offers the strongest practical trade-off.
Finally, to assess the impact of fine-tuning on code-styled data, we evaluate Llama-3.1-8B-Instruct on HumanEval (Chen et al., 2021) before and after fine-tuning. We find that its Pass@1 score increases slightly from $72.6$ to $73.1$, indicating that fine-tuning for the code style pattern does not degrade the model’s coding capability. As noted in §5.1, the fine-tuning data used for the code style pattern contains code-styled instruction–response pairs rather than actual programming content. Our goal is to simulate superficial style alignment, not to introduce new code knowledge. Therefore, a large improvement in coding ability is not expected.
@@
[…]
Defense List Poem News Legal Shakespeare Code
Strategy Llama Qwen Llama Qwen Llama Qwen Llama Qwen Llama Qwen Llama Qwen
No Defense +0.264+0.264 +0.259+0.259 +0.357+0.357 +0.395+0.395 +0.398+0.398 +0.405+0.405 +0.589+0.589 +0.630+0.630 +0.162+0.162 +0.154+0.154 +0.360+0.360 +0.395+0.395
vanilla 0.076-0.076 $\bm{-0.034}$ +0.208+0.208 +0.213+0.213 +0.080+0.080 +0.087+0.087 +0.096+0.096 +0.066+0.066 0.016-0.016 0.021-0.021 0.028-0.028 0.014-0.014
PTST +0.010+0.010 +0.005+0.005 +0.072+0.072 +0.144+0.144 +0.149+0.149 +0.177+0.177 +0.359+0.359 +0.381+0.381 0.247-0.247 0.241-0.241 0.034-0.034 0.020-0.020
SPPFT +0.170+0.170 +0.113+0.113 +0.212+0.212 +0.265+0.265 +0.243+0.243 +0.278+0.278 +0.309+0.309 +0.306+0.306 0.073-0.073 0.135-0.135 +0.244+0.244 +0.259+0.259
Constrained +0.001+0.001 0.002-0.002 0.002-0.002 0.001-0.001 0.017-0.017 0.015-0.015 0.085-0.085 0.048-0.048 0.003-0.003 +0.005+0.005 +0.006+0.006 +0.006+0.006
Paraphrase 0.062-0.062 0.029-0.029 +0.296+0.296 +0.285+0.285 +0.107+0.107 +0.113+0.113 +0.095+0.095 +0.066+0.066 0.019-0.019 0.054-0.054 0.032-0.032 0.014-0.014
SafeStyle $\bm{-0.081}$ $\bm{-0.034}$ $\bm{-0.027}$ $\bm{-0.019}$ $\bm{-0.083}$ $\bm{-0.088}$ $\bm{-0.116}$ $\bm{-0.076}$ $\bm{-0.408}$ $\bm{-0.362}$ $\bm{-0.092}$ $\bm{-0.065}$
[…]
Defense News Legal
Strategy $\DeltaASR Δ\DeltaLC$_WR $\DeltaASR Δ\DeltaLC$_WR
No Defense +0.063+0.063 +24.260+24.260 +0.013+0.013 +27.840+27.840
Vanilla +0.017+0.017 +25.667+25.667 +0.001+0.001 +28.215+28.215
PTST +0.040+0.040 $\bm{+25.840}$ +0.005+0.005 +21.484+21.484
SPPFT +0.074+0.074 +23.797+23.797 +0.012+0.012 +23.920+23.920
Constrained 0.001-0.001 +21.068+21.068 $\bm{-0.004}$ +22.606+22.606
Paraphrase +0.015+0.015 +24.929+24.929 0.001-0.001 +28.627+28.627
SafeStyle $\bm{-0.007}$ +25.137+25.137 $\bm{-0.004}$ $\bm{+28.865}$
Table 9: Safety–utility ranking based on Tables 1 and 2, where 11 denotes the best rank. SafeStyle achieves the strongest overall safety–utility trade-off.
Defense Safety Utility Average
Strategy Ranking Ranking Ranking
No Defense $6.458$ $3.542$ $5.000$
Vanilla $3.583$ $3.375$ $3.479$
PTST $4.000$ $5.000$ $4.500$
SPPFT $6.042$ $3.875$ $4.958$
Constrained $3.542$ $5.042$ $4.292$
Paraphrase $3.333$ $\bm{3.333}$ $3.333$
SafeStyle $\bm{1.042}$ $3.833$ $\bm{2.438}$
[…]
Defense News Legal
Strategy $\Delta\textbf{ASR}_{\textbf{ori}}$ $\Delta\textbf{ASR}_{\textbf{ave}}$ $\Delta\textbf{LC\_WR}_{\textbf{ori}}$ $\Delta\textbf{LC\_WR}_{\textbf{ave}}$ $\Delta\textbf{ASR}_{\textbf{ori}}$ $\Delta\textbf{ASR}_{\textbf{ave}}$ $\Delta\textbf{LC\_WR}_{\textbf{ori}}$ $\Delta\textbf{LC\_WR}_{\textbf{ave}}$
No Defense +0.008+0.008 +0.016+0.016 +7.385+7.385 +5.559+5.559 +0.523+0.523 +0.563+0.563 +1.032+1.032 +3.766+3.766
Vanilla 0.003-0.003 0.010-0.010 +2.003+2.003 +3.992+3.992 0.005-0.005 0.009-0.009 +2.255+2.255 +4.411+4.411
PTST 0.002-0.002 0.006-0.006 +6.802+6.802 +5.958+5.958 +0.001+0.001 0.009-0.009 +4.548+4.548 +4.749+4.749
SPPFT +0.003+0.003 +0.010+0.010 $\bm{+7.249}$ $\bm{+5.967}$ +0.011+0.011 +0.005+0.005 +4.403+4.403 +4.440+4.440
Constrained 0.001-0.001 0.010-0.010 +4.670+4.670 +4.010+4.010 0.006-0.006 0.013-0.013 +4.462+4.462 +4.418+4.418
Paraphrase 0.004-0.004 0.007-0.007 +1.548+1.548 +4.191+4.191 0.005-0.005 0.008-0.008 +2.663+2.663 +4.616+4.616
SafeStyle $\bm{-0.006}$ $\bm{-0.016}$ +7.092+7.092 +5.888+5.888 $\bm{-0.009}$ $\bm{-0.015}$ $\bm{+4.815}$ $\bm{+4.832}$